OpenStudy (anonymous):

pleaaaaaase help ill fan/medal/bake you cookies/give you the soul of my first born

OpenStudy (anonymous):

Please explain this step by step I'm so lost

OpenStudy (campbell_st):

ok.... part 1... to find the 1st 4 terms you have \[-4 \times (\frac{1}{3})^{n - 1}\] 1st term substitute n = 1 and evaluate -4 2nd term substitute n = 2 and evaluate -4/3 3rd term substitute n = 3 and evaluate -4/9 I'll let you do the 4th term (b) the series converges since the common ratio, r = 1/3. whenever a geometric series where the the common ratio r -1 < r < 1 the series converges. (c) to find the sum to infinity of infinite sum use \[S_{\infty} = \frac{a}{1 - r}\] a is the 1st term.... you found it in part a here is a good site with information about geometric series http://www.mathsisfun.com/algebra/sequences-sums-geometric.html

OpenStudy (anonymous):

Ok I'm kind of understanding now